OpenStudy (selfie_flawless):

Which of these equations represents the commutative property of addition? (a + b) + c = c + (a + b) (a + b) + c = (a + b) + c (a + b) + c = a + b + c (a + b) + c = a + (b + c)

OpenStudy (danjs):

The commutative property says it doesn't matter which order you decide to add things together. For example if you have A and B. the commutative property shows you can add them in either order A + B = B + A

OpenStudy (selfie_flawless):

but what is the answer?

OpenStudy (danjs):

the one that is structured like that.. order doesnt matter (a + b) + c = c + (a + b)